Direct Materials and Direct Labor Variances At the beginning of June, Kimber Toy Company budgeted 26,000 toy action figures to be manufactured in June at standard direct materials and direct labor costs as follows: Direct materials $36,400 Direct labor 15,600 Total $52,000 The standard materials price is $0.40 per pound. The standard direct labor rate is $15.00 per hour. At the end of June, the actual direct materials and direct labor costs were as follows: Actual direct materials $32,600 Actual direct labor 14,000 Total $46,600 There were no direct materials price or direct labor rate variances for June. In addition, assume no changes in the direct materials inventory balances in June. Kimber Toy Company actually produced 22,600 units during June. Determine the direct materials quantity and direct labor time variances. Round your per unit computations to two decimal places, if required. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. Direct materials quantity variance $ Direct labor time variance.
Answer #1

a) Direct material quantity variance = (Standard cost-actual cost) = (36400/26000*22600)-32600 = 960 U

Direct labor time variance = (Standard cost-actual cost) = (15600/26000*22600)-14000 = 440 U